# Durham, UK Mamba $\times$ Transformer for Image Inpainting

### v2

@@ Abstract
Image inpainting, or image completion, is a crucial task in computer vision that aims to restore missing or damaged regions of images with semantically coherent content. This technique requires a precise balance of local texture replication and global contextual understanding to ensure the restored image integrates seamlessly with its surroundings. Traditional methods using Convolutional Neural Networks (CNNs) are effective at capturing local patterns but often struggle with broader contextual relationships due to the limited receptive fields. Recent advancements have incorporated transformers, leveraging their ability to understand global interactions. However, these methods face computational inefficiencies and struggle to maintain fine-grained details. To overcome these challenges, we introduce M×$\times$×T composed of the proposed Hybrid Module (HM), which combines Mamba with the transformer in a synergistic manner. Mamba is adept at efficiently processing long sequences with linear computational costs, making it an ideal complement to the transformer for handling long-scale data interactions. Our HM facilitates dual-level interaction learning at both pixel and patch levels, greatly enhancing the model to reconstruct images with high quality and contextual accuracy. We evaluate M×$\times$×T on the widely-used CelebA-HQ and Places2-standard datasets, where it consistently outperformed existing state-of-the-art methods.
@@ 1 Introduction
Image inpainting, as known as image completion, aims at restoring missing or damaged parts of images with semantically plausible context. This demands accurate modeling of both global and local information within the corrupted image, which is crucial as the global and local interaction maintain the coherence of both the content and style of the missing areas, ensuring seamless integration with the surrounding image regions [Yu et al.(2018)Yu, Lin, Yang, Shen, Lu, and Huang].
@@
In this paper, we introduce M×$\times$×T, consisted by proposed Hybrid Modules that synergistically combine the strengths of both transformer and Mamba. This novel approach allows for a dual-level interaction learning from the patch level and pixel level. Our comparative experiments demonstrate that our M×$\times$×T outperforms existing state-of-the-art methods on two wildly used datasets, CelebA-HQ and Places2. We summarize our contributions as: 1) We propose M×$\times$×T to introduce Mamba combined with transformer for image inpainting. 2) We design a novel Hybrid Module to capture the feature interaction in both pixel-level and patch-level. 3) Our M×$\times$×T overall suppress the state-of-the-art methods on both CelebA-HQ and Places2 datasets. 4) M×$\times$×T is able to adapt to high-resolution images with only training on low-resolution data.
@@ 2.1 Image Inpainting
Image inpainting is an ill-posed low-level vision task that aims to infer the missing regions of the image via the undamaged pixels. Conventional works employ diffusion-based approaches to inpaint the missing regions by deriving by neighbouring visible pixels [Sridevi and Srinivas Kumar(2019)], or filling the missing areas by using the good match patches from the background or external sources such as the depth information [Atapour-Abarghouei et al.(2016)Atapour-Abarghouei, de La Garanderie, and Breckon, Barnes et al.(2009)Barnes, Shechtman, Finkelstein, and Goldman]. Although these methods can effectively complete small missing regions, they face challenges in precisely reconstructing more complex scenes due to limited global understanding of the image. Recently, deep learning based image inpainting studies propose to use CNN-based encoder-decoder architecture [Yan et al.(2018)Yan, Li, Li, Zuo, and Shan, Yu et al.(2020)Yu, Guo, Jin, Wu, Chen, Li, Zhang, and Liu, Suin et al.(2021)Suin, Purohit, and Rajagopalan] or CNN-based Generative Adversarial Networks (GANs) [Xu et al.(2014)Xu, Ren, Liu, and Jia, Yu et al.(2018)Yu, Lin, Yang, Shen, Lu, and Huang, Iizuka et al.(2017)Iizuka, Simo-Serra, and Ishikawa, Yi et al.(2020)Yi, Tang, Azizi, Jang, and Xu, Peng et al.(2021)Peng, Liu, Xu, and Li, Sargsyan et al.(2023)Sargsyan, Navasardyan, Xu, and Shi], which significantly improves the visual plausibility and diversity of inpainted images. However, the limited convolutional receptive field hinders the model’s learning of long-range dependencies, which motivates studies on expanding the receptive field by applying the frequency convolution techniques [Suvorov et al.(2022)Suvorov, Logacheva, Mashikhin, Remizova, Ashukha, Silvestrov, Kong, Goka, Park, and Lempitsky, Chu et al.(2023)Chu, Chen, Sun, Lian, Wang, Zuo, Zhao, Xing, and Lu] or developing transformer-based models [Chen et al.(2024)Chen, Atapour-Abarghouei, and Shum, Li et al.(2022a)Li, Lin, Zhou, Qi, Wang, and Jia, Chen et al.(2021)Chen, Wang, Guo, Xu, Deng, Liu, Ma, Xu, Xu, and Gao, Liang et al.(2021)Liang, Cao, Sun, Zhang, Van Gool, and Timofte]. Nonetheless, practical training and application of the transformer-based models are still constrained by the exponential complexity of self-attention calculations. In particular, it is still challenging to use pixel-level self-attention to achieve relatively high-resolution image inpainting. To this end, we turn our eyes to the field of selective SSM (i.e., Mamba [Gu and Dao(2023)]) for achieving image long-range pixel-wise dependency learning with robust spatial awareness.
@@ 2.2 SSMs in Computer Vision
Recently, State Space Models (SSMs) have demonstrated promising advantages of long sequence modeling and linear-time complexity in Natural Language Processing (NLP) [Gu et al.(2021)Gu, Goel, and Re]. This work specifically tackles the problem of vanishing gradients in SSMs when solving the exponential function by the linear first-order Ordinary Differential Equations [Gu et al.(2020)Gu, Dao, Ermon, Rudra, and Ré]. Building on the rigorous theoretical proofs of the HiPPO framework that enables SSMs to capture long-range dependencies, Gu et al. [Gu and Dao(2023)] further introduce a data-dependent selective structure SSM (i.e., Mamba) to significantly improve the computational efficiencies in conventional SSMs. Inspired by pioneering SSMs, vision-specific adaptations of the Mamba architecture, such as Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and V-Mamba [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], propose visual SSMs designs for computer vision tasks including image classification and object detection [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. However, their performance are still behind of the state-of-the-art transformer-based models like SpectFormer [Patro et al.(2023)Patro, Namboodiri, and Agneeswaran], SVT [Patro and Agneeswaran(2024)], and WaveViT [Yao et al.(2022)Yao, Pan, Li, Ngo, and Mei]. U-Mamba [Ma et al.(2024)Ma, Li, and Wang] effectively extends the capabilities of Mamba for biomedical image segmentation by proposing a hybrid CNN-SSM block. However, these studies ineffectively leverage the capabilities of Mamba in image long-range pixel-level dependency learning and overlook the critical spatial awareness during model designs.
@@ 3.1 State Space Models and Mamba
The State Space Model (SSM) is generally known as a linear time-invariant system that map a 1-dimensional input sequence $x(t)\in\mathbb{R}$ to a response $y(t)\in\mathbb{R}$ via a hidden latent state $h(t)\in\mathbb{R}^{N}$ (Eq.1). For efficient linear-complexity deep learning model training, the structured SSMs employ a zero-order hold discretization rule (Eq.2) to transform the continues parameter $(\Delta,\textbf{{A}},\textbf{{B}})$ to discrete parameters $(\overline{\textbf{{A}}},\overline{\textbf{{B}}})$, as shown in Eq.3.
@@
where $\textbf{{A}}\in\mathbb{R}^{N\times N}$, $\textbf{{B}}\in\mathbb{R}^{N\times 1}$, $\textbf{{C}}\in\mathbb{R}^{1\times N}$, and $\Delta$ is a time-scale parameter. Mamba [Gu and Dao(2023)], one of the most recent selective SSMs, introduces a gated selective mechanism to propagate or eliminate selected information based on the current state, significantly improves the content-reasoning performance. Specifically, Mamba changes the model from time-invariant to time-varying via converting parameters $\Delta$, B, C into input-dependent functions.
@@ 4 Method
The overall pipeline of the proposed M×$\times$×T is illustrated in Fig. 1, which is a U-Net shape architecture formed with 7 Hybrid Blocks. Formally, the masked image ${I}_{masked}\in\mathbb{R}^{H\times W\times 3}$ concatenated with a mask ${M}\in\mathbb{R}^{H\times W\times 1}$ as the input ${I}_{in}$. We first use an overlapped convolution to embed ${I}_{in}$, then feed into the following 7 HBs with 3 times downsampling and 3 times upsampling. At the end, one convolution layer projects the final output ${I}_{out}$. Each Hybrid Block is consisted of $n$ Hybrid Modules (HM), as shown in Fig. 1 (b), where $n$ is the number of HMs. Each HM has a Transformer block, a Mamba block and a Context Broadcasting Feedforward Network (CBFN), which will be detailed in the section 4.1.
@@ 4.1 Hybrid Module
where $Att$ is the attention map. In this work, the spatial dimension is reduced to $8$. After multiplying $Att$ and $V^{\prime}$, we get the initial output $F^{\prime\prime}$. To further enhance local context, we incorporate a Local Enhancement operation $LE(V)$ as proposed in [Ren et al.(2022)Ren, Zhou, He, Feng, and Wang], which is implemented using a $3\times 3$ depth-wise convolution, to effectively balances capturing extensive global interactions with detailed local features. After an element-wise addition, resulting the final output of SRSA:
@@ 5 Experiment Results
Implementation Details. All experiments are carried out on one Nvidia A100 GPU. During training, we adopt the Adam optimiser [Kingma et al.(2020)Kingma, Ba, and Adam] with ${\beta}_{1}$ = 0.9, ${\beta}_{2}$ = 0.999. The learning rate is set to $1\times 10^{-4}$ and the batch size is 4. In the Hybrid Blocks, the number of Hybrid Modules are [4,6,6,8,6,6,4].
@@ 5.1 Comparison with State of the Art
Quantitative Comparison. For a fair comparison, we employ the officially released models and test them with the same test sets and masks. As shown in table 1, our M×$\times$×T output perform in all metrics cross different mask ratios. Especially on CelebA-HQ, at the increasing mask ratios, M×$\times$×T improves PSNR by 2.0%, 2.3% and 1.8% respectively, and decreases LPIPS by 12.3%, 11.6% and 10.5% respectively. Moreover, for the Places2, our model demonstrated comparable performance to SOTAs such as MISF and LAMA. While our training utilized the Places2-Standard dataset with 1.8 million images, MISF and LAMA were trained on the Places2-Challenge dataset, which contains 8 million images. Despite using only 22.5% of the images employed by these benchmarks, our model achieved comparable results, showcasing its robustness and efficiency.
@@ 5.3 Application: High-Resolution Image Inpainting
Our M×$\times$×T is designed with linear computational complexity, enabling it to effectively handle high-resolution image inpainting task. We directly apply our model, pre-trained on Places2-standard dataset, to real-world high-resolution images, demonstrating its capability. The example is illustrated in Fig. 3.
@@ 5.4 Limitation and Discussion
Future Work. In this work, our goal is to develop an inpainting model capable of reconstructing high-quality images, emphasizing fine-grained details and contextually plausible structures. Moving forward, our next objective is to enhance its capability by integrating multimodal foundational models, such as CLIP, to make the inpainting process controllable through text guidance, allowing users to influence the reconstruction results with descriptive input, while maintaining the high quality.

### v3

# Durham, UK Corresponding Author: Hubert P. H. Shum Mamba $\times$ Transformer for Image Inpainting

@@ Abstract
Image inpainting, or image completion, is a crucial task in computer vision that aims to restore missing or damaged regions of images with semantically coherent content. This technique requires a precise balance of local texture replication and global contextual understanding to ensure the restored image integrates seamlessly with its surroundings. Traditional methods using Convolutional Neural Networks (CNNs) are effective at capturing local patterns but often struggle with broader contextual relationships due to the limited receptive fields. Recent advancements have incorporated transformers, leveraging their ability to understand global interactions. However, these methods face computational inefficiencies and struggle to maintain fine-grained details. To overcome these challenges, we introduce M×$\times$×T composed of the proposed Hybrid Module (HM), which combines Mamba with the transformer in a synergistic manner. Mamba is adept at efficiently processing long sequences with linear computational costs, making it an ideal complement to the transformer for handling long-scale data interactions. Our HM facilitates dual-level interaction learning at both pixel and patch levels, greatly enhancing the model to reconstruct images with high quality and contextual accuracy. We evaluate M×$\times$×T on the widely-used CelebA-HQ and Places2-standard datasets, where it consistently outperformed existing state-of-the-art methods. The code will be released: https://github.com/ChrisChen1023/MxT.
@@ 1 Introduction
Image inpainting, as known as image completion, aims at restoring missing or damaged parts of images with semantically plausible context. This demands accurate modeling of both global and local information within the corrupted image, which is crucial as the global and local interaction maintains the coherence of both the content and style of the missing areas, ensuring seamless integration with the surrounding image regions [Yu et al.(2018)Yu, Lin, Yang, Shen, Lu, and Huang].
@@
In this paper, we introduce M×$\times$×T, consisting of proposed Hybrid Modules that synergistically combine the strengths of both transformer and Mamba. This novel approach allows for dual-level interaction learning from the patch level and pixel level. Our comparative experiments demonstrate that our M×$\times$×T outperforms existing state-of-the-art methods on two wildly used datasets, CelebA-HQ and Places2. We summarize our contributions as: 1) We propose M×$\times$×T to introduce Mamba combined with transformer for image inpainting. 2) We design a novel Hybrid Module to capture the feature interaction at both the pixel level and patch level. 3) Our M×$\times$×T overall suppress the state-of-the-art methods on both CelebA-HQ and Places2 datasets. 4) M×$\times$×T is able to adapt to high-resolution images with only training on low-resolution data.
@@ 2.1 Image Inpainting
Image inpainting is an ill-posed low-level vision task that aims to infer the missing regions of the image via the undamaged pixels. Conventional works employ diffusion-based approaches to inpaint the missing regions by deriving by neighbouring visible pixels [Sridevi and Srinivas Kumar(2019)], or filling the missing areas by using the good match patches from the background or external sources such as the depth information [Atapour-Abarghouei et al.(2016)Atapour-Abarghouei, de La Garanderie, and Breckon, Barnes et al.(2009)Barnes, Shechtman, Finkelstein, and Goldman]. Although these methods can effectively complete small missing regions, they face challenges in precisely reconstructing more complex scenes due to limited global understanding of the image. Recently, deep learning based image inpainting studies propose to use CNN-based encoder-decoder architecture [Yan et al.(2018)Yan, Li, Li, Zuo, and Shan, Yu et al.(2020)Yu, Guo, Jin, Wu, Chen, Li, Zhang, and Liu, Suin et al.(2021)Suin, Purohit, and Rajagopalan] or CNN-based Generative Adversarial Networks (GANs) [Xu et al.(2014)Xu, Ren, Liu, and Jia, Yu et al.(2018)Yu, Lin, Yang, Shen, Lu, and Huang, Iizuka et al.(2017)Iizuka, Simo-Serra, and Ishikawa, Yi et al.(2020)Yi, Tang, Azizi, Jang, and Xu, Peng et al.(2021)Peng, Liu, Xu, and Li, Sargsyan et al.(2023)Sargsyan, Navasardyan, Xu, and Shi], which significantly improves the visual plausibility and diversity of inpainted images. However, the limited convolutional receptive field hinders the model’s learning of long-range dependencies, which motivates studies on expanding the receptive field by applying the frequency convolution techniques [Suvorov et al.(2022)Suvorov, Logacheva, Mashikhin, Remizova, Ashukha, Silvestrov, Kong, Goka, Park, and Lempitsky, Chu et al.(2023)Chu, Chen, Sun, Lian, Wang, Zuo, Zhao, Xing, and Lu] or developing transformer-based models [Chen et al.(2024)Chen, Atapour-Abarghouei, and Shum, Li et al.(2022a)Li, Lin, Zhou, Qi, Wang, and Jia, Chen et al.(2021)Chen, Wang, Guo, Xu, Deng, Liu, Ma, Xu, Xu, and Gao, Liang et al.(2021)Liang, Cao, Sun, Zhang, Van Gool, and Timofte, Chang et al.(2023)Chang, Koulieris, and Shum]. Nonetheless, practical training and application of the transformer-based models are still constrained by the exponential complexity of self-attention calculations. In particular, it is still challenging to use pixel-level self-attention to achieve relatively high-resolution image inpainting. To this end, we turn our eyes to the field of selective SSM (i.e., Mamba [Gu and Dao(2023)]) for achieving image long-range pixel-wise dependency learning with robust spatial awareness.
@@ 2.2 SSMs in Computer Vision
Recently, State Space Models (SSMs) have demonstrated promising advantages of long sequence modeling and linear-time complexity in Natural Language Processing (NLP) [Gu et al.(2021)Gu, Goel, and Re]. This work specifically tackles the problem of vanishing gradients in SSMs when solving the exponential function by the linear first-order Ordinary Differential Equations [Gu et al.(2020)Gu, Dao, Ermon, Rudra, and Ré]. Building on the rigorous theoretical proofs of the HiPPO framework that enables SSMs to capture long-range dependencies, Gu et al. [Gu and Dao(2023)] further introduce a data-dependent selective structure SSM (i.e., Mamba) to significantly improve the computational efficiencies in conventional SSMs. Inspired by pioneering SSMs, vision-specific adaptations of the Mamba architecture, such as Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and V-Mamba [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], propose visual SSMs designs for computer vision tasks including image classification and object detection [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. However, their performance is still behind of the state-of-the-art transformer-based models like SpectFormer [Patro et al.(2023)Patro, Namboodiri, and Agneeswaran], SVT [Patro and Agneeswaran(2024)], and WaveViT [Yao et al.(2022)Yao, Pan, Li, Ngo, and Mei]. U-Mamba [Ma et al.(2024)Ma, Li, and Wang] effectively extends the capabilities of Mamba for biomedical image segmentation by proposing a hybrid CNN-SSM block. However, these studies ineffectively leverage the capabilities of Mamba in image long-range pixel-level dependency learning and overlook the critical spatial awareness during model designs.
@@ 3.1 State Space Models and Mamba
The State Space Model (SSM) is generally known as a linear time-invariant system that maps a 1-dimensional input sequence $x(t)\in\mathbb{R}$ to a response $y(t)\in\mathbb{R}$ via a hidden latent state $h(t)\in\mathbb{R}^{N}$ (Eq.1). For efficient linear-complexity deep learning model training, the structured SSMs employ a zero-order hold discretization rule (Eq.2) to transform the continues parameter $(\Delta,\textbf{{A}},\textbf{{B}})$ to discrete parameters $(\overline{\textbf{{A}}},\overline{\textbf{{B}}})$, as shown in Eq.3.
@@
where $\textbf{{A}}\in\mathbb{R}^{N\times N}$, $\textbf{{B}}\in\mathbb{R}^{N\times 1}$, $\textbf{{C}}\in\mathbb{R}^{1\times N}$, and $\Delta$ is a time-scale parameter. Mamba [Gu and Dao(2023)], one of the most recent selective SSMs, introduces a gated selective mechanism to propagate or eliminate selected information based on the current state, significantly improving the content-reasoning performance. Specifically, Mamba changes the model from time-invariant to time-varying via converting parameters $\Delta$, B, C into input-dependent functions.
@@ 4 Method
The overall pipeline of the proposed M×$\times$×T is illustrated in Fig. 1, which is a U-Net shape architecture formed with 7 Hybrid Blocks. Formally, the masked image ${I}_{masked}\in\mathbb{R}^{H\times W\times 3}$ concatenated with a mask ${M}\in\mathbb{R}^{H\times W\times 1}$ as the input ${I}_{in}$. We first use an overlapped convolution to embed ${I}_{in}$, then feed into the following 7 HBs with 3 times downsampling and 3 times upsampling. At the end, one convolution layer projects the final output ${I}_{out}$. Each Hybrid Block consists of $n$ Hybrid Modules (HM), as shown in Fig. 1 (b), where $n$ is the number of HMs. Each HM has a Transformer block, a Mamba block and a Context Broadcasting Feedforward Network (CBFN), which will be detailed in section 4.1.
@@ 4.1 Hybrid Module
where $Att$ is the attention map. In this work, the spatial dimension is reduced to $8$. After multiplying $Att$ and $V^{\prime}$, we get the initial output $F^{\prime\prime}$. To further enhance local context, we incorporate a Local Enhancement operation $LE(V)$ as proposed in [Ren et al.(2022)Ren, Zhou, He, Feng, and Wang], which is implemented using a $3\times 3$ depth-wise convolution, to effectively balances capturing extensive global interactions with detailed local features. After an element-wise addition, the final output of SRSA is:
@@ 5 Experiment Results
Implementation Details. All experiments are carried out on one Nvidia A100 GPU. During training, we adopt the Adam optimiser [Kingma et al.(2020)Kingma, Ba, and Adam] with ${\beta}_{1}$ = 0.9, ${\beta}_{2}$ = 0.999. The learning rate is set to $1\times 10^{-4}$ and the batch size is 4. In the Hybrid Blocks, the numbers of Hybrid Modules are [4,6,6,8,6,6,4].
@@ 5.1 Comparison with State of the Art
Quantitative Comparison. For a fair comparison, we employ the officially released models and test them with the same test sets and masks. As shown in table 1, our M×$\times$×T outperforms in all metrics across different mask ratios. Especially on CelebA-HQ, at the increasing mask ratios, M×$\times$×T improves PSNR by 2.0%, 2.3% and 1.8% respectively, and decreases LPIPS by 12.3%, 11.6% and 10.5% respectively. Moreover, for Places2, our model demonstrated comparable performance to SOTAs such as MISF and LAMA. While our training utilized the Places2-Standard dataset with 1.8 million images, MISF and LAMA were trained on the Places2-Challenge dataset, which contains 8 million images. Despite using only 22.5% of the images employed by these benchmarks, our model achieved comparable results, showcasing its robustness and efficiency.
@@ 5.3 Application: High-Resolution Image Inpainting
Our M×$\times$×T is designed with linear computational complexity, enabling it to effectively handle high-resolution image inpainting tasks. We directly apply our model, pre-trained on the Places2-standard dataset, to real-world high-resolution images, demonstrating its capability. The example is illustrated in Fig. 3.
@@ 5.4 Limitation and Discussion
Future Work. In this work, our goal is to develop an inpainting model capable of reconstructing high-quality images, emphasizing fine-grained details and contextually plausible structures. Moving forward, our next objective is to enhance its capability by integrating multimodal foundational models, such as CLIP, to make the inpainting process controllable through text guidance, allowing users to influence the reconstruction results with descriptive input, while maintaining high quality.
@@ 6 Conclusion
7 Acknowledgement
This research is supported in part by the EPSRC NortHFutures project (ref: EP/X031012/1).